# Data-Driven Goal Recognition Design for General Behavioral Agents

### v2

@@ 5.2.2 Experiment 2: Evaluating goal recognition design
We next evaluate the performance of our approach that incorporates the data-driven model of human behavior from experiment 1. To do this, we randomly generate 30 initial environments using the same setup as in our simulations. These environments are then updated according to four different methods, all operating within a modification budget of 20.
@@
We recruited 200 workers from MTurk. Each worker was randomly assigned to one of the four treatments above, with the distinction between treatments being the environments presented to them. Workers were randomly assigned a goal for each environment and tasked with navigating to reach it. We utilize the collected data to evaluate different goal recognition design approaches.
@@
Figure 7: Comparing different goal recognition design approaches in our human subject experiments. Our approach coupled with data-driven models is shown to generate environments that enable the most effective goal recognition with real human decision makers.
@@
Comparing empirical overlapping actions. To evaluate the effectiveness of each goal recognition design approach, we first measure the empirical overlapping actions towards each of the two goals. Specifically, each recruited worker is exposed to all 30 environments in their assigned treatment and is instructed to reach one of two randomly selected goals. For each treatment, we calculate the number of overlapping actions for every pair of workers assigned to different goals within each environment. Therefore, for each treatment, we obtain a distribution of the number of overlapping actions. This metric reflects the difficulty of inferring agent goals for the environment and serves as an empirical proxy for wcd. Figure 7(a) presents the percentiles of the number of overlapping actions for each treatment. In particular, in the lowest 25%, the environments by our approaches have no overlapping actions, indicating that the environment makes it easy to infer the goal of the agents from observations. Generally, our approach, when integrated with data-driven models, leads to environments with fewer number of overlapping actions and therefore facilitate the easier identification of the humans’ intended goal.
@@
Comparing the accuracy of goal inference. Instead of using wcd or proxy of wcd for evaluations, we next directly measure whether we can indeed infer the human goals based on partial observations of their actions, utilizing an off-the-shelf Bayesian inference algorithm. [^3] We assume an initial uniform prior distribution. As the agent takes actions, we update the posterior belief over goals based on the likelihood probability of the data-driven behavior model prediction. Specifically, for each worker, we reveal the first $k$ portion of the worker’s actions to the inference algorithm to infer the worker’s goal. This enables us to compute the average inference accuracy. The results in Figure 7(b) demonstrate that our approach leads to environments that are easier for goal recognition.
6 Conclusion
Effective human-AI collaboration hinges on the AI’s ability to infer the goals of humans. In this study, we work on the problem of goal recognition design, updating the decision-making environments to make it easier for the AI agents to perform goal recognition. By addressing the key limitations in the existing literature, notably the computational demand and the assumption of optimal agent behavior, we have developed a data-driven optimization framework that is both efficient and scalable. Through simulations and human-subject experiments, we show that our approach outperforms state-of-the-art approaches in standard settings, applies to settings that existing approaches cannot address, and leads to environments easier for goal recognition with real-world human decision makers.
Societal impact. Our approach is capable of integrating general models of various agent behavior, making it more applicable to real-world scenarios involving human agents. However, with more accurate human models, it also opens up the concerns of privacy and potential mis-use of the techniques for unethical surveillance. While our work does not directly facilitate such misuse, the broader field of goal recognition research could potentially be exploited.
Limitations and future work. While our work has extended the literature in goal recognition design to address suboptimal agent behavior, more flexible budget constraints, and more complicated environments, there are still several limitations. In particular, we have assumed static and full observability of the environment. Extending our framework to deal with more complex environments, particularly those that are dynamic and partially observable would be an interesting and important future work. Moreover, we have assumed that user behavior remains static, so we can learn a human model from their historical behavior. Finally, given our framework is adaptable to different forms of agent behavior, understanding how different patterns of agent suboptimality impacts the difficulty of goal recognition could lead to useful insights for enabling better human-AI collaboration.
Acknowledgements
@@
This work is supported in part by J.P. Morgan Faculty Research Award and a Global Incubator Seed Grant from McDonnell International Scholars Academy.
@@ Appendix A Experiment Details and Additional Results
In this section, we provide details of the experiments that have not been included in the main paper due to space constraints. We also include and discuss additional experiment results.
@@ A.1 Implementation Details
In this subsection, we discuss the details of our implementation for training the predictive model for wcd. To train a CNN-based prediction model, we generated an extensive training dataset by solving wcd for a large number of randomly selected environments, given a specific agent model. For creating a predictive model for environments with a grid size of $N$ by $N$, we utilized a custom Convolutional Neural Network (CNN) architecture. This architecture is tailored to process input data, $w_{i}$, consisting of $k$ $N\times N$ channels, where $k$ represents the number of potential objects in the environment.
The architecture comprises two distinct blocks: the initial block contains three successive convolutional layers, each followed by batch normalization and a rectified linear unit (ReLU) activation function. The subsequent block introduces pooling operations and additional convolutional layers for more sophisticated feature extraction. After the convolutional blocks, the architecture includes fully connected layers, leading to a single output unit. The design integrates Leaky ReLU, ReLU, and dropout layers to add non-linearity and provide regularization. This flexible architecture can adapt to tasks with multi-channel input data, allowing for modifications based on the specific needs of the task and dataset characteristics. Batch normalization and dropout are incorporated to promote training stability and prevent overfitting. We employed the Adam optimizer and Mean Squared Error (MSE) loss for training the models, specifically focusing on settings involving optimal human agents.
To create the dataset to train the model, we randomly generated $150$K grid designs and computed the corresponding wcd. This dataset was then split into the training dataset ($80\%$) and validation dataset ($20\%$). For our experiments reported in Section 5 with $13\times 13$ grid size and optimal agent behavior, we have achieved an MSE value around 0.18 for both training and validation errors. The results are shown in Figure 8. We have also examined a variety of setups, and generally we can reach small errors for the predictive model that can enable the optimization.
Figure 8: Training loss and validation errors for the predictive model for wcd with $13\times 13$ grid world and optimal human behavior.
A.2 Results for Run Time Comparisons
@@
Figure 9: Run time taken by each method in the different experimental conditions. In the standard setup with a small grid size, all methods except exhaustive complete within less than half of a second but only our approach scales with a larger grid size and other more complex configurations.
All our experiments are conducted on a cluster of 40 CPU cores (Intel Xeon Gold 6148 CPU @@@@2.40GHz), 1 GPU (NVIDIA Tesla V100 SXM2 32GB), and a maximal memory of 80GB.
@@ Appendix B Additional Experiment Results
In this section, we report the additional experiment results that are not included in the main text due to space constraints. In Section 5.1.2 of the main paper, we provided details of our performance with a large grid size. In a smaller grid world, our approach achieves comparable performance to the baselines but it significantly outperforms them in larger grid sizes as shown in Figure 10.
@@ Appendix C Details of Simulations
To generate our training and evaluation datasets, we randomly generated environments and kept environments that are valid, e.g., the goals are reachable, and that the objects don’t overlap. Below we provide more details for environment generation for specified grid sizes for both overcooked and grid-world domains.
@@ C.1 Gridworld
In a grid world, a valid environment includes a starting position, blocked cells, and two goal positions. The starting position is randomly placed in the first column, and the goal positions are randomly placed in the last two columns. The number of blocked positions in each grid is randomly selected from a range of 00 to 2×2$\times$2 × the grid width. We discard any assignments that make the goals unreachable. For experiments with suboptimal behavior, we also randomly assigned small subgoal rewards to unblocked cells that the agent would collect on its way to the goal state. The two goal states assigned a large reward that is 10 times the largest subgoal reward.
@@ C.2 Overcooked-AI
In Overcooked-AI, a valid environment includes one pot, one tomato source, one onion source, one dish source, one serving point, no open spaces at the border, and any number of open spaces and blocked cells. All objects must be reachable from the agent’s randomly assigned starting position, with the number and positions of blocked cells assigned randomly. The agent is randomly assigned any two possible goals: three tomato soups, three onion soups, or a mixed soup. Each goal has the same randomly assigned reward value. Suboptimal behavior is modeled by assigning small random subgoal rewards when adding ingredients to the pot.

### v3

@@ 5.2.2 Experiment 2: Evaluating Goal Recognition Design
We next evaluate our approach that incorporates the data-driven model of human behavior from Experiment 1. We randomly generate 30 initial environments using the same setup as in simulations. These environments are then updated according to four different methods, all operating within a modification budget of 20.
@@
We recruited 200 workers from Amazon Mechanical Turk. Each worker was randomly assigned to one of the four treatments above, with the distinction between treatments being the environments presented to them. Workers were randomly assigned a goal for each environment and tasked with navigating to reach it. We utilize the collected data to evaluate GRD design approaches.
@@
Figure 6: Comparing different GRD approaches in our human subject experiments. Our approach coupled with data-driven models is shown to generate environments that enable the most effective goal recognition with real human decision-makers.
@@
Comparing empirical overlapping actions. To evaluate the effectiveness of each GRD approach, we first measure the empirical overlapping actions toward each of the two goals. Specifically, each recruited worker is exposed to all 30 environments in their assigned treatment and is instructed to reach one of two randomly selected goals. For each treatment, we compute the number of overlapping actions for every pair of workers assigned to different goals within the same environment. This yields a distribution of overlapping actions for each treatment. This metric reflects the difficulty of inferring an agent’s goal and serves as an empirical proxy for wcd. Figure 6(a) presents the percentiles of the number of overlapping actions across treatments. Notably, in the lowest 25%, environments generated by our approach show no overlapping actions, indicating that goal inference is relatively easy in these settings. Overall, when combined with data-driven models, our approach produces environments with statistically significantly fewer overlapping actions (paired t-test, p-value < 0.005), thereby facilitating more accurate identification of human goals.
@@
Comparing the accuracy of goal inference. Instead of evaluating the effectiveness of our approaches based on wcd or its proxy, we next directly assess whether human goals can be accurately inferred from partial observations of their actions. To do this, we use an off-the-shelf Bayesian inference algorithm (Murphy, 2012)  [^3] We assume a uniform prior distribution over goals. As the agent takes action, we update the posterior belief over goals through Bayesian updates based on the likelihood probability of the data-driven behavior model prediction.. As the agent acts, we update the posterior belief using the likelihood derived from our data-driven behavior model. Specifically, for each worker, we reveal the first $k$ portion of their actions to the inference algorithm, which then attempts to infer the worker’s goal. This allows us to compute the average inference accuracy. The results, shown in Figure 6(b), demonstrate that our approach produces environments that make goal recognition easier.
6 Conclusion and Future Work
Effective human-AI collaboration hinges on the AI system’s ability to infer human goals. In this work, we introduce a data-driven framework for goal recognition design (GRD) that optimizes environments to enhance goal disambiguation. Our method addresses two major limitations of prior approaches: computational inefficiency and the unrealistic assumption of fully optimal agent behavior. At the core of our framework is a learned oracle that predicts goal recognition difficulty, worst-case distinctiveness (wcd). By integrating this oracle into a gradient-based optimization process, we avoid expensive exact evaluations and unlock a scalable, flexible design procedure. This enables three core advantages: (1) improved scalability, (2) support for general, potentially suboptimal agent models, and (3) extensibility to diverse constraints and objectives.
Our empirical evaluation highlights the effectiveness and versatility of the approach. In both classical grid worlds and complex Overcooked-AI environments, our method consistently outperforms baseline techniques in reducing wcd while significantly lowering computational cost. Crucially, it remains effective under suboptimal agent behavior—demonstrating robustness where many existing methods break down. Human-subject experiments further validate the practical utility of our framework, showing that it generalizes beyond simulation and maintains its advantage even when agent models are learned from real human data. To our knowledge, this is the first work to evaluate GRD methods in a human-subject setting. Through ablation studies, we demonstrate that the predictive oracle plays a critical role in guiding effective environment optimization, with more accurate or expressive models leading to significantly better outcomes.
Limitations and future work. While our framework advances goal recognition design, several limitations remain. First, our approach requires a large amount of data to train accurate predictive oracles. Although simulations can be used to generate training data, data from novel and richer environments is often scarce and presents challenges for simulation. Future work could explore several methods to address this limitation, including leveraging transfer learning (Pan & Yang, 2009) by pretraining on simulated environments and fine-tuning on limited data from new environments, as well as using generative models to produce synthetic yet behaviorally plausible training data (Goodfellow et al., 2014; Xu et al., 2019). Active learning techniques that reduce the amount of data required during collection also offer a promising direction. While these strategies require further validation, we believe they represent potentially viable and interesting avenues for applying GRD in data-scarce settings.
Second, we assume fully observable environments and stationary human behavior in this work. To address these limitations, future work could train the predictive model on inputs derived from belief state representations, incorporating uncertainty in agent observations (Kaelbling et al., 1998) into the training process. In addition, the predictive model could be periodically fine-tuned using newly accumulated behavioral data, allowing it to adapt to changes in human decision-making over time.
Moreover, our investigation focuses on single-agent, discrete environments. To extend our framework to multi-agent settings, future work would need to model agent behavior in the presence of other agents, accounting for their interactions. To support continuous environments, the Lagrangian relaxation can be adapted to penalize changes using continuous distance functions such as norms, enabling gradient-based optimization without altering the underlying pipeline. Overall, the modularity of our framework enables future work to address these limitations in a targeted and flexible manner.
Broader impact. Finally, while our work aims to advance human–AI collaboration through goal recognition design, it is important to acknowledge the societal risks that may arise from its misuse.
First, designing environments that make human goals and intentions easier to infer could be misused. Systems capable of anticipating intentions might be deployed not to assist users but to manipulate or steer their behavior. Because system designers typically hold an advantageous position over decision makers within these environments, such capabilities can further amplify their influence over individuals. Second, since GRD relies on predictive models trained on human behavioral data, it raises privacy concerns. Even when raw data are not directly exposed, adversaries could analyze model outputs or behaviors to infer sensitive personal attributes, creating opportunities for profiling or surveillance. Third, by making goal inference more accurate and efficient, GRD could foster excessive reliance on AI predictions. Such overreliance risks undermining human autonomy and weakening critical judgment, especially if users begin to accept AI inferences uncritically. Taken together, these risks highlight how a technique intended to enhance collaboration could, under certain deployments, compromise user welfare and agency.
In light of these concerns, we discuss possible avenues for risk mitigation. On the technical side, privacy-preserving methods such as differential privacy could be integrated into the training of behavioral models, limiting the extent to which sensitive information is encoded and thereby reducing opportunities for exploitation. Moreover, incorporating uncertainty-aware methods into GRD could ensure that systems explicitly communicate the confidence of their inferences. Such mechanisms would help mitigate both privacy risks and overreliance by reminding users of the fallibility of AI systems. On the policy side, once the capabilities and risks of GRD are more fully understood, it will be important to weigh the potential benefits against the possible harms. This evaluation should guide the development of regulatory frameworks and ethical guidelines—for example, requiring that GRD-driven modifications not significantly reduce user autonomy relative to standard environments, and mandating transparency about the objectives and data underlying deployment. Establishing such safeguards will be essential to ensure that GRD is developed and applied responsibly, and that its benefits are realized without compromising broader societal values.
Acknowledgments
@@
This work is supported in part by a J.P. Morgan Faculty Research Award and a Global Incubator Seed Grant from the McDonnell International Scholars Academy.
@@ Appendix A Experiment Details and Additional Results
In this section, we provide details of the experiments not included in the main paper due to space constraints. We also include and discuss additional experimental results.
@@ A.1 Implementation Details
This subsection describes the implementation of the predictive model. We trained the predictive models on a simulated dataset generated by solving wcd for randomly sampled environments with a specific agent behavior model. Several architectures were tested, including CNNs, Linear Models, Kernel Ridge Regression, and Transformers, with CNNs performing the best in predicting wcd. The models processes inputs of the shape $k\times N\times N$ inputs, where $k$ is the number of potential objects in an $N\times N$ grid.
Figure 7: Training loss and validation errors for the CNN predictive model for wcd with $13\times 13$ grid world and optimal human behavior.
The CNN architecture used across the paper employs a ResNet18 backbone with an adapted $k$-channel input layer, followed by three fully connected layers. Outputs are reduced to 32 and 16 dimensions with Leaky ReLU activations and dropout, culminating in a single non-negative scalar via ReLU activation.
For the experiments in Section 5, we trained a CNN-based model on 100K data points, split into 80% training and 20% validation. The model achieved an MSE below 0.18 for both sets, with the validation set used for hyperparameter tuning. Figure 7 shows the training loss for the wcd predictive model for the grid world domain ($13\times 13$) assuming optimal behavior. We used Adam optimizer and MSE loss and tested learning rates of 0.1, 0.01, 0.001, and 0.0001. A learning rate of 0.001 consistently produced the lowest validation error, enabling reliable wcd reduction across different setups.
A.2 More Results for Run Time Comparisons
@@
Figure 8: Run time taken by each method in the different experimental conditions. In the standard setup with a small grid size, all methods except exhaustive complete within less than half of a second but our approach scales better with more complex configurations.
@@ Appendix B Additional Experiment Results
In this section, we report the additional experiment results that are not included in the main text due to space constraints. In Section 4 of the main paper, we provided details of our performance with a large grid size. In a smaller grid world, our approach achieves comparable performance to the baselines but it significantly outperforms them in larger grid sizes as shown in Figure 9.
@@ B.1 Definition of Budget
B.2 Langrange Parameter Space
In the main paper, we described how we relax the original constrained optimization problem by introducing a Lagrangian penalty parameter $\lambda$, which indirectly controls the budget through penalizing constraint violations. Since the budget cannot be specified directly, we instead vary $\lambda$ across a predefined range to explore different trade-offs between optimizing the objective and satisfying the budget constraint. Specifically, we sweep over a log-scaled set of values:, $\lambda\in\{0,0.001,0.002,0.005,0.007,0.01,0.02,0.05,0.07,0.1,0.2,0.5,0.7,1.0,2,5,7\}$. This range covers several orders of magnitude, allowing the optimization to balance constraint enforcement from very loose (small $\lambda$) to very strict (large $\lambda$) penalties. For all experiments, including those with a single budget constraint, we used this same set of $\lambda$ values. In scenarios involving two constraints, such as Blocking and Unblocking constraints, we applied this range independently to each constraint, effectively exploring the joint penalty space by combining these $\lambda$ values for each constraint.
B.3 Ablation Analysis
In the main paper, we provided the ablation analysis for the optimal behavior model with a grid size of 13. Table 4 provides the same analysis for grid size 6. The results show mixed performance in this smaller environment, and our approach sometimes underperforms baselines. This reflects that our method’s advantages emerge primarily in larger, more complex environments where traditional approaches become computationally intractable.
[…]
Predictive Model Loss (mse) Optimization Approach wcd Reduction
Train Validation Standard Individual Shared
[…]
@@ Appendix C Details of Simulations
To generate our training and evaluation datasets, we randomly generated environments and kept valid environments, e.g., those in which the goals are reachable and the objects don’t overlap. Below, we provide more details about environment generation for specified grid sizes for both overcooked and grid-world domains.
@@ C.1 Grid world
In a grid world, a valid environment includes a starting position, blocked cells, and two goal positions. The starting position is randomly placed in the first column, and the goal positions are randomly placed in the last two columns. The number of blocked positions in each grid is randomly selected from a range of 0 to $2\times$ the grid width. We discard any assignments that make the goals unreachable. For experiments with suboptimal behavior, we also randomly assigned small subgoal rewards to unblocked cells that the agent would collect on its way to the goal state. The two goal states are assigned a large reward that is 10 times the largest subgoal reward.
@@ C.2 Overcooked-AI
In Overcooked-AI, a valid environment includes one pot, one tomato source, one onion source, one dish source, one serving point, no open spaces at the border, and any number of open spaces and blocked cells. All objects must be reachable from the agent’s randomly assigned starting position, with the number and positions of blocked cells assigned randomly. The agent is randomly assigned one of two goals for each experiment, where both goals are randomly selected from the set of three possible recipes: three-tomato soup, three-onion soup, or mixed soup. Each goal has the same randomly assigned reward value. Suboptimal behavior is modeled by assigning small random subgoal rewards when adding ingredients to the pot.
@@ Appendix D Details of Human-Subject Experiment
Appendix E Computation Details
All experiments were run on a computing cluster equipped with 40 CPU cores (Intel Xeon Gold 6148 @ 2.40GHz), a single NVIDIA Tesla V100 SXM2 GPU (32GB), and up to 80GB of memory. To ensure reproducibility, we worked within a clean and consistent software environment built around Python 3.10 and widely used scientific libraries. PyTorch was our main deep learning framework, with NumPy and pandas handling numerical computation and data processing. During model training and the discrete gradient descent optimization, we relied on PyTorch’s autograd system to compute gradients automatically and efficiently.